# Deep Reinforcement Learning Enhanced Rate-Splitting Multiple Access for Interference Mitigation

### v1

@@ I Introduction
Rate-splitting multiple access (RSMA) effectively manages interference in many scenarios, such as in ultra-dense 6G networks [7] or in satellite and aerial integrated networks [8]. It aims to attain a balance between treating interference as noise and fully decoding unintended messages. To achieve this, RSMA splits messages into common and private layers and uses successive interference cancellation (SIC) at the receivers to enable a step-by-step decoding process for all common messages and for individual private messages. Ensuring receivers decode common messages before extracting private information, rate-splitting offers a new strategy to optimize communication in interference. Since its introduction, rate-splitting has been a focal point in information theory, shaping advancements in interference management within modern communication networks [9]. Researchers explore its applications, complexities, performance boundaries, and practical implementations, highlighting its significance [10] in enhancing communication efficiency in high-interference environments.
@@
While RSMA offers high flexibility, tackling the non-convex optimization challenges for RSMA is particularly challenging [11]. This difficulty intensifies in scenarios with multiple users or increased rate-splitting layers, creating a complex web of potential configurations. Consequently, finding the best rate-splitting combination becomes computationally challenging, necessitating sophisticated algorithms and robust computational resources, such as MATLAB CVX tools [12].
@@
Deep reinforcement learning (DRL) is pivotal for addressing these challenges in RSMA [13]. RSMA’s complexity requires advanced navigation in high-dimensional spaces, and DRL, within a Markov decision process (MDP), refines rate-splitting in dynamic interference scenarios, enhancing communication efficiency. Reinforcement learning excels in handling uncertainties, optimizing resource allocation, and precoder optimization [14]. Furthermore, multi-agent deep reinforcement learning (MADRL) employs decentralized decision-making and aligns with distributed communication systems, effectively managing non-stationary and non-linear dynamics in channels.
@@ I-A Related Literature
Research on interference channels has been pivotal in information theory and communications, with seminal contributions shaping our understanding. Han and Kobayashi’s study in 1981 [15] proposed rate-splitting as a unique method for interference channels, introducing it as an innovative technique. Rate-splitting divides transmitted messages into common and private layers, addressing interference challenges by balancing noise treatment and efficient decoding [7].
@@
[19]
MADQN, MADDQN, MAD3QN
Broadcast channel
Pilot contamination
Cell-free massive MIMO
[20, 21]
MADDPG/DDPG
Interference channel
Precoder
Multi-user Multi-cell, single-cell MISO
[22, 23]
PPO
Broadcast channel
Precoder and power allocation coefficient
Multi-user single-cell SISO
[24]
TRPO
Broadcast channel
Energy harvesting, sensing and communication capabilities
Multi-user single-cell MISO
[25]
DQL
Interference channel
Beamformer and power
Multi-user multi-cell MISO
[26]
DRL
Interference channel
Beamformer and power
Single-user multi-cell MISO
[27]
DQL
Broadcast channel
Energy efficiency
Multi-user single-cell MISO
[28]
MADQL
Broadcast channel
Energy efficiency and power
Multi-user single-cell MISO
[29, 30, 9]
DL/DDPG
Broadcast channel
Precoder and power
Multi-user single-cell MIMO
This work
MADDPG
Interference channel
Precoder and power allocation coefficient
multi-user SISO, multi-user MISO, multi-user MIMO
@@ II System Model
To satisfy the power constraints at each one of the transmitters, we assume $|{\bm{w}}_{ikc}|^{2}+|{\bm{w}}_{ikp}|^{2}\leq P_{ik}$, where $\sum_{k=1}^{M_{i}}P_{ik}=P_{i}$ is the total power of $BS_{i}$, and $\bm{w}_{ikc}$ and $\bm{w}_{ikp}$ represent the $k$th column of $\bm{W}_{ic}$ and $\bm{W}_{ip}$ respectively, $k=1,\ldots,Q_{i}$  [^1] The values the variable $k$ takes depend on the value $i$. However, to keep the notation simple, we slightly abuse the notation and do not have an $i$ index for $k$.. Moreover, we define
@@
$\bm{P}_{ip}$ $=$ $\left[P_{i1p},\ldots,P_{iQ_{i}p}\right]^{T}=\left[|\bm{w}_{i1p}|^{2},\ldots,|\bm{w}_{iQ_{i}p}|^{2}\right]^{T},$ (4)
@@
In order to be able to write the achievable rates with RSMA, the decoding order for $\bm{b}_{ic}$ and $\bm{b}_{ip}$ have to be determined at both users [32]. Moreover, to attain the best possible achievable rates, one has to consider all possible choices of these decoding orders. For the particular system model we study, we take 2 different decoding orders into consideration for each one of the users. Namely, ${UE}_{i}$ either decodes in the order (a) or (b) in (6).
@@
$\max_{\bm{W}_{ic},\bm{W}_{ip},\bm{P}_{ic},\bm{P}_{ip},i=1,2}$ $\qquad\beta R_{1}+(1-\beta)R_{2}$ (16a)
s.t. $\qquad|\bm{w}_{ikc}|^{2}+|\bm{w}_{ikp}|^{2}\leq P_{ik},$ (16b)
$\quad\quad\quad\quad\quad i=1,2,\quad k=1,2,..,Q_{i}$
$\qquad\quad\sum_{k=1}^{M_{i}}P_{ik}=P_{i},\quad i=1,2.$ (16c)
@@ III MADDPG for Precoding and Power Allocation Coefficients Optimization
Figure 1: System architecture for MADDPG with RSMA for MIMO interference channels.
@@
Figure 2: MADDPG with RSMA algorithm structure for the MIMO case.
First, we assume a fixed decoding order. For each decoding order, we consider an environment with two agents. The overall system architecture is shown in Fig. 1 and the algorithm structure is summarized in Fig. 2. In this system, each agent $i$ at $BS_{i}$ chooses a precoding matrix $\bm{W}_{i}=[\bm{W}_{ic}~{}\bm{W}_{ip}]$ and a power allocation coefficient vector $\bm{P}_{ic}$ based on the local observation $O_{i}=[\bm{H}_{i}~{}\bm{G}_{i}]$, $i\in\{1,2\}$. We construct actor networks parameterized by $\phi_{i1}$ for the precoding vectors $\bm{w}_{ik}$, for $k=1,\ldots,Q_{i}$; and parameterized by $\phi_{i2}$ for the power allocation coefficients $P_{ikc}$, and a critic network to evaluate the performance of policies for each agent $i$. Please note that, for the single antenna case, actor network is only responsible for choosing $P_{ikc}$ for each agent since there is no need to determine $\bm{W}_{i}$. Let $\mu=\{\mu_{\phi_{11}}(O_{1}),\mu_{\phi_{12}}(O_{1}),\mu_{\phi_{21}}(O_{2}),\mu_{\phi_{22}}(O_{2})\}$ denote the set of policies parameterized by $\phi=\{\phi_{1},\phi_{2}\}$ where $\phi_{1}=\{\phi_{11},\phi_{12}\}$ and $\phi_{2}=\{\phi_{21},\phi_{22}\}$. The agents will choose their actions $a_{i}=[P_{i1c}~{}P_{i2c}~{}\cdots~{}P_{iQ_{i}c}~{}\bm{w}_{i1c}~{}\bm{w}_{i1p}~{}\bm{w}_{i2c}~{}\bm{w}_{i2p}~{}\cdots~{}\bm{w}_{iQ_{i}c}~{}\bm{w}_{iQ_{i}p}]$ according to the partial state $O_{i}$ by following a deterministic policy $a_{i}=\mu_{\phi_{i}}(O_{i})$. To ensure sufficient exploration, we also add a noise vector, whose entries are independent and identically distributed according to $\mathcal{N}(0,\,\sigma_{N}^{2})$ to the deterministic action $a_{i}=\mu_{\phi_{i}}(O_{i})$. Then, the gradient of the expected reward $J(\phi_{i})$ for each agent $i$ can be computed as
$\nabla_{\phi_{i}}J(\phi_{i})=\mathbb{E}\left[\nabla_{a_{i}}\mathcal{Q}_{\theta_{i}}^{\mu}(s,a_{1},a_{2})|_{a_{i}=\mu_{\phi_{i}}(O_{i})}\nabla_{\phi_{i}}\mu_{\phi_{i}}(O_{i})\right],$ (17)
@@ III-A Decoding Order Optimization
         Compute target Q-value:
@@ III-B Imperfect Channel State Information
where the resulting estimation error terms, denoted by $\bm{E}_{i}$ and $\hat{\bm{E}}_{j}$, arise from the disparity between the estimated channel coefficients, represented by $\tilde{\bm{H}}_{i}$ and $\tilde{\bm{G}}_{j}$, and the actual channel coefficients denoted as $\bm{H}_{i}$ and $\tilde{\bm{G}}_{j}$. It is important to note that the matrices $\bm{E}_{i}$, $\bm{H}_{i}$, and $\tilde{\bm{H}}_{i}$ all belong to the complex field and have dimensions $N_{i}\times M_{i}$; i.e. $\mathbb{C}^{N_{i}\times M_{i}}$. In the case of varying imperfections, each entry $[\bm{E}_{i}]{(x,y)}$, $x=1,\ldots,N_{i}$, $y=1,\ldots,M_{i}$, is determined as $\frac{SNR^{-0.6}}{5}\mathcal{CN}(0,\sigma^{2})$, where $\sigma^{2}=1$. Similarly, the matrices $\hat{\bm{E}}_{j}$, $\bm{G}_{j}$, and $\tilde{\bm{G}}_{j}$ are all in $\mathbb{C}^{N_{i}\times M_{j}}$, $i\neq j$. Each entry in $\hat{\bm{E}}_{i}$ is drawn from $\frac{SNR^{-0.6}}{5}\mathcal{CN}(0,\sigma^{2})$ with $\sigma^{2}=1$. In the case of fixed channel estimation errors, each entry in $\bm{E}_{i}$ or $\hat{\bm{E}}_{j}$ will be distributed according to $\frac{SNR^{-0.6}}{5}\mathcal{CN}(0,\sigma^{2})$ with $\sigma^{2}=1$ in Section V.

### v2

@@ I Introduction
Among these sophisticated techniques, rate-splitting is among the most robust and effective to mitigate interference. This method was first proposed in 1981 [7]. For example, in a two-user interference channel, there are two transmitters and two receivers, both of whom are only interested in their corresponding messages. In an interference channel, the regular approach is to treat the other user’s message as noise. However, this other user’s message is not noise, it has a structure and can be decoded if desired. To create a balance between treating unwanted messages as noise and fully decoding them, other user’s messages can be partially decoded, and the remaining part can be treated as noise. This balance is achieved by constructing the messages in common and private parts. Rate-splitting uses successive interference cancellation (SIC) at the receivers to enable a step-by-step decoding process for all common messages and for individual private messages. This adaptability makes the rate-splitting method unique and achieves the largest rate region in interference channels. Later, it was discovered that multi-antenna broadcast channels require as much interference mitigation as interference channels, and rate-splitting is quite useful in ultra-dense 6G networks [8], in integrated satellite and aerial networks [9] or in reconfigurable intelligent surface-aided communication systems [10]. Since its introduction, rate-splitting has been a focal point in information theory, shaping advances in interference management. Researchers explore its applications, complexities, performance boundaries, and practical implementations, highlighting its significance in enhancing communication efficiency in high-interference environments in both interference and multi-antenna broadcast channels [11].
@@
While RSMA offers high flexibility, tackling the non-convex optimization challenges for RSMA is particularly challenging [12]. Algorithms like the iterative weighted minimum mean square error (WMMSE) algorithm [13] can be utilized to address the traditional non-convex sum-rate maximization problem. However, they come with significant computational complexity due to the need for high-dimensional matrix inversions and an excessive number of iterations. This challenge becomes more severe in scenarios with multiple users, unknown or complex channel characteristics, or increased rate-splitting layers. Consequently, finding the best rate-splitting combination becomes computationally challenging, necessitating sophisticated algorithms and robust computational resources, such as MATLAB CVX tools [14].
Another practical approach is to use lower-complexity heuristic methods, such as zero forcing (ZF) or maximum ratio transmission (MRT) [15]. Although these algorithms are computationally efficient, their performance is often limited because ZF disregards residual interference, while MRT focuses solely on maximizing signal strength without considering interference management or user-specific channel conditions. As a result, they do not achieve satisfactory system performance in our scenario.
@@
Deep reinforcement learning (DRL) is pivotal for addressing these challenges in RSMA [16]. RSMA’s complexity requires advanced navigation in high-dimensional spaces, and DRL, within a Markov decision process (MDP), refines rate-splitting in dynamic interference scenarios, enhancing communication efficiency. Reinforcement learning excels in handling uncertainties, optimizing resource allocation, and precoder optimization [17]. In terms of computational complexity, the main advantage of DRL method is that it delegates the majority of the complex computations to the training phase, leaving only simple forward calculations to be performed during the implementation phase. Unlike standard deep learning (DL) algorithms, DRL algorithms can adapt to changing channel conditions and does not require a fixed, large dataset for training. Furthermore, multi-agent deep reinforcement learning (MADRL) employs decentralized decision-making and aligns with distributed communication systems such as interference channels considered in this paper, and effectively manages the present nonlinear dynamics.
@@ I-A Related Literature
Unlike the recent literature on RSMA, which is for multi-antenna broadcast channels, in this paper we consider rate-splitting for the interference channel. Research on interference channels has been pivotal in information theory and communications, with seminal contributions shaping our understanding. Han and Kobayashi’s study in 1981 [7] proposed rate-splitting as a unique method for interference channels, introducing it as an innovative technique. Rate-splitting divides transmitted messages into common and private layers, addressing interference challenges by balancing noise treatment and efficient decoding [8].
@@
[21] MADQN, MADDQN, MAD3QN Broadcast channel Pilot contamination Cell-free massive MIMO
[22, 23] MADDPG/DDPG Interference channel Precoder Multi-user Multi-cell, single-cell MISO
[24, 25] PPO Broadcast channel Precoder and power allocation coefficient Multi-user single-cell SISO
[26] TRPO Broadcast channel Energy harvesting, sensing and communication capabilities Multi-user single-cell MISO
[27] DQL Interference channel Beamformer and power Multi-user multi-cell MISO
[28] DRL Interference channel Beamformer and power Single-user multi-cell MISO
[29] DQL Broadcast channel Energy efficiency Multi-user single-cell MISO
[30] MADQL Broadcast channel Energy efficiency and power Multi-user single-cell MISO
[31, 32, 10] DL/DDPG Broadcast channel Precoder and power ✓/✗ Multi-user single-cell MIMO
This work MADDPG Interference channel Precoder and power allocation coefficient multi-user SISO, multi-user MISO, multi-user MIMO
@@ II System Model
To satisfy the power constraints at each one of the transmitters, we assume $|{\bm{w}}_{ikc}|^{2}+|{\bm{w}}_{ikp}|^{2}\leq P_{ik}$, where $\sum_{k=1}^{Q_{i}}P_{ik}=P_{i}$ is the total power of $BS_{i}$, and $\bm{w}_{ikc}$ and $\bm{w}_{ikp}$ represent the $k$th column of $\bm{W}_{ic}$ and $\bm{W}_{ip}$ respectively, $k=1,\ldots,Q_{i}$  [^1] The values the variable $k$ takes depend on the value $i$. However, to keep the notation simple, we slightly abuse the notation and do not have an $i$ index for $k$.. Moreover, we define
@@
$\bm{P}_{ip}$ $=$ $\left[P_{i1p},\ldots,P_{iQ_{i}p}\right]^{T}=\left[|\bm{w}_{i1p}|^{2},\ldots,|\bm{w}_{iQ_{i}p}|^{2}\right]^{T}$ (4)
@@
In order to be able to write the achievable rates with RSMA, the decoding order for $\bm{b}_{ic}$ and $\bm{b}_{ip}$ must be determined in both users [34]. Moreover, to attain the best possible achievable rates, one has to consider all possible choices of these decoding orders. For the particular system model we study, we take 2 different decoding orders into consideration for each one of the users. Namely, ${UE}_{i}$ either decodes in the order (a) or (b) in (6).
@@
$\max_{\bm{W}_{ic},\bm{W}_{ip},~{}i=1,2}$ $\qquad\beta R_{1}+(1-\beta)R_{2}$ (16a)
s.t. $\qquad\sum_{k=1}^{Q_{i}}\left(|\bm{w}_{ikc}|^{2}+|\bm{w}_{ikp}|^{2}\right)\leq P_{i},\quad i=1,2.$ (16b)
@@ III MADDPG for Precoding and Power Allocation Coefficients Optimization
[…]
Figure 1: System and MADDPG with RSMA algorithm structure for MIMO interference channels. In the MADDPG algorithm, there is centralized training and distributed execution.
Initially, we assume the optimal decoding order that maximizes the sum-rate and without loss of generality, we consider an environment with two agents each representing a user. The overall system architecture and the algorithm structure are summarized in Fig. 1. Later, in Section III-A, we will extend this approach by determining the optimal decoding order using MADDP.
In this system, each agent $i$ at $BS_{i}$ chooses a precoding matrix $\bm{W}_{i}=[\bm{W}_{ic}~{}\bm{W}_{ip}]$ and a power allocation coefficient vector $\bm{P}_{ic}$ based on the local observation $O_{i}=[\bm{H}_{i}~{}\bm{G}_{i}]$, $i\in\{1,2\}$. We construct actor networks parameterized by $\phi_{i1}$ for the precoding matrix $\bm{W}_{i}=[\bm{W}_{ic}~{}\bm{W}_{ip}]$; and parameterized by $\phi_{i2}$ for the power allocation coefficients $\bm{P}_{ic}$, and a critic network to evaluate the performance of policies for each agent $i$. Please note that, in the single antenna case, the actor network is solely responsible for selecting $\bm{P}_{ic}$. Here, ${P}_{ic}$ becomes a single power allocation coefficient for each agent, eliminating the need to determine $\bm{W}_{i}$. Let $\mu=\{\mu_{\phi_{11}}(O_{1}),\mu_{\phi_{12}}(O_{1}),\mu_{\phi_{21}}(O_{2}),\mu_{\phi_{22}}(O_{2})\}$ denote the set of policies parameterized by $\phi=\{\phi_{1},\phi_{2}\}$ where $\phi_{1}=\{\phi_{11},\phi_{12}\}$ and $\phi_{2}=\{\phi_{21},\phi_{22}\}$. The agents will choose their actions $a_{i}=[P_{i1c}~{}P_{i2c}~{}\cdots~{}P_{iQ_{i}c}~{}\bm{w}_{i1c}~{}\bm{w}_{i1p}~{}\bm{w}_{i2c}~{}\bm{w}_{i2p}~{}\cdots~{}\bm{w}_{iQ_{i}c}~{}\bm{w}_{iQ_{i}p}]$ according to the partial state $O_{i}$ by following a deterministic policy $a_{i}=\mu_{\phi_{i}}(O_{i})$. To ensure sufficient exploration, we also add a noise vector, whose entries are independent and identically distributed according to $\mathcal{N}(0,\,\sigma_{N}^{2})$ to the deterministic action $a_{i}=\mu_{\phi_{i}}(O_{i})$. Then, the gradient of the expected reward $J(\phi_{i})$ for each agent $i$ can be computed as
$\nabla_{\phi_{i}}J(\phi_{i})$
$=\mathbb{E}\left[\nabla_{a_{i}}\mathcal{Q}_{\theta_{i}}^{\mu}(s,a_{1},a_{2})|_{a_{i}=\mu_{\phi_{i}}(O_{i})}\nabla_{\phi_{i}}\mu_{\phi_{i}}(O_{i})\right],$ (17)
@@ III-A Decoding Order Optimization
         Compute target $\mathcal{Q}$-value:
@@ III-B Imperfect Channel State Information
where the resulting estimation error terms, denoted by $\bm{E}_{i}$ and $\hat{\bm{E}}_{j}$, arise from the disparity between the estimated channel coefficients, represented by $\tilde{\bm{H}}_{i}$ and $\tilde{\bm{G}}_{j}$, and the actual channel coefficients denoted as $\bm{H}_{i}$ and $\tilde{\bm{G}}_{j}$. It is important to note that the matrices $\bm{E}_{i}$, $\bm{H}_{i}$, and $\tilde{\bm{H}}_{i}$ all belong to the complex field and have dimensions $N_{i}\times M_{i}$; i.e. $\mathbb{C}^{N_{i}\times M_{i}}$. For the second case, where estimation errors decrease with increasing SNR, each entry $[\bm{E}_{i}]{(x,y)}$, $x=1,\ldots,N_{i}$, $y=1,\ldots,M_{i}$, is determined as $\frac{SNR^{-0.6}}{5}\mathcal{CN}(0,\sigma^{2})$, where $\sigma^{2}=1$. Similarly, the matrices $\hat{\bm{E}}_{j}$, $\bm{G}_{j}$, and $\tilde{\bm{G}}_{j}$ are all in $\mathbb{C}^{N_{i}\times M_{j}}$, $i\neq j$. Each entry in $\hat{\bm{E}}_{i}$ is drawn from $\frac{SNR^{-0.6}}{5}\mathcal{CN}(0,\sigma^{2})$ with $\sigma^{2}=1$. In the case of fixed channel estimation errors, each entry in $\bm{E}_{i}$ or $\hat{\bm{E}}_{j}$ will be distributed according to $\frac{10^{-0.6}}{5}\mathcal{CN}(0,\sigma^{2})$ with $\sigma^{2}=1$ in Section V.
@@ III-C MADDPG with no Rate-Splitting
III-D Complexity Analysis of the Proposed Algorithm
Training in MADDPG is the phase where agents interact with their environment, improve their decision-making policies, and adjust the parameters of the actor and critic networks. This phase is the most computationally expensive because it requires a lot of calculations to learn the best policies and value functions. Key steps in training involve several important processes. First, each agent stores its experiences in a replay buffer, which includes states, actions, and rewards from previous time steps. Then, each agent updates its policy by computing gradients based on feedback from the critic network, which estimates the value of state-action pairs. The critic network itself updates its value estimates using a loss function that incorporates the Bellman equation. Finally, agents must balance exploration and exploitation, requiring multiple interactions with the environment to explore new strategies while exploiting known ones. Specifically, the training complexity can be expressed as
$O\left(E\left(N_{\text{agents}}KQ_{a}^{2}+L^{2}MKB^{2}\right)+TD\right.$ (26)
$\quad\quad\quad\quad\quad\quad\quad+N_{\text{agents}}\left.\left(KQ_{a}^{2}+L^{2}MKB^{2}\right)\right)$
where $E$ is the number of training episodes or time steps, $N_{\text{agents}}$ is the number of agents, $K$ is the action space size, $Q_{a}$ is the size of the $\mathcal{Q}$-value function, $L$ is the number of layers in the network, $M$ is the number of units(hidden size) in each layer, $B$ is the batch size, $T$ is the episode length, and $D$ is the replay memory length.
The expression (26) involves several key components. The forward pass complexity is $O(N_{\text{agents}}KQ_{a}^{2})$ for each agent, which is required to calculate $\mathcal{Q}$-values and process them through the network. The network layers’ complexity is $O(L^{2}MKB^{2})$, accounting for the complexities of the neural network layers in both the actor and the critic networks. The experience replay complexity is $O(TD)$, representing the effort required to store and sample experiences from the replay buffer during training. The critic update complexity is $O(N_{\text{agents}}(KQ_{a}^{2}+L^{2}MKB^{2}))$, which includes the complexity of calculating the Bellman equation and performing back-propagation for the critic network. Together, these components provide a comprehensive picture of the computational complexity of MADDPG, covering the training process, experience replay, and critic updates, all scaled by the number of agents and training steps. The forward calculation complexity is low because, once the model is trained, the forward pass becomes computationally efficient. It simply involves a feed forward operation through the actor and critic networks to select actions and estimate values. Since there are no gradients to compute or weights to update during the forward pass, the process is much simpler. Essentially, the forward pass only requires matrix multiplications for each network, which is far less computationally demanding compared to the training phase. This is a key reason why execution complexity is low across most deep learning algorithms once the model has been trained.
In MADDPG, after training, the model is deployed in a distributed setting where each agent executes its own policy in parallel, independently interacting with the environment and selecting actions based on the trained model. The key steps in distributed execution include parallel execution, where multiple agents execute their policies concurrently, each having its own local environment; action selection, where each agent uses the trained actor model to predict actions based on its local state; environment interaction, where agents interact with the environment and obtain feedback (reward and next state); and state update, where each agent updates its local state in real time based on the environment’s response.
The complexity of distributed execution can be expressed as
$O(N_{\text{agents}}(Q_{a}+Q_{c}))$ (27)
where $N_{\text{agents}}$ is the number of agents in the distributed system, and $Q_{a}$ and $Q_{c}$ are the output sizes of the actor and critic networks for action and value predictions. The distributed execution complexity is low due to several reasons. First, parallelization allows each agent to independently perform a forward pass to select its action, with no interactions between agents during this phase, except for sharing the environment. Second, no training is involved in this phase, as there is no back propagation or gradient update; agents only compute actions based on the trained model. Lastly, real-time execution makes the complexity low per agent, and the load can be efficiently scaled by distributing tasks across agents.
Complexity of simple precoders like ZF is primarily determined by the need to solve matrix inversion problems, which has a computational cost of $O(N_{i}^{3}+N_{i}^{2}K)$, where $N_{i}$ is the number of antennas and $K$ is the number of users. Advanced RSMA schemes, which require convex optimization, scale to $O(N_{i}^{3}K^{3})$. However, this does not account for the added complexity of decoding order selection, which can be substantial, especially when exhaustive search methods are used. This can add $O(K!)$ to the overall complexity, though practical implementations often employ heuristic methods that reduce this to $O(K^{2})$. When compared with MADDPG, which operates in a distributed manner, the complexity is quite different. MADDPG’s execution complexity scales to (27) represent the complexities of each agent’s action and critic evaluation, respectively. This distributed execution of MADDPG results in a significantly lower overall complexity compared to traditional RSMA optimization, such as weighted MMSE [13], especially as the number of users or agents increases, demonstrating the efficiency of MADDPG in multi-agent settings relative to the centralized nature of RSMA.